OpenStudy (love_to_love_you):

If A and B are matrices of the same order, then cA + cB is __equal/not equal to___ c(B + A) and cA − cB is __equal/not equal to___ c(B − A).

OpenStudy (love_to_love_you):

I get confused when it comes to matrices

hartnn (hartnn):

you can factor out a constant from matrices. pX+pY = p (X+Y) pX-pY = p (X-Y)

OpenStudy (love_to_love_you):

So they're both equal?

hartnn (hartnn):

1st one is equal. are you sure about 2nd?

OpenStudy (love_to_love_you):

I didn't think it was equal for second but got confused

hartnn (hartnn):

in that link, look for 'Properties of Scalar Multiplication' cA-cB = c(A-B) and A-B is not same as B-A

OpenStudy (love_to_love_you):

So it isn't equal to?

hartnn (hartnn):

yes, the 2nd one is "not equal to"

OpenStudy (love_to_love_you):

Alrighty, thank you :)

OpenStudy (love_to_love_you):

And thank you for the link @Bobo-i-bo

OpenStudy (bobo-i-bo):

One way we can think about it is like this: \[cA-cB=cA+(-c)B=cA+(c(-1))B=cA+c((-1)B)=cA+c(-B)=c(A-B)\]

OpenStudy (bobo-i-bo):

Np ^_^